OpenStudy (anonymous):

Give a possible formula for the function below. Let A = (-1, 4) and B = (1, 1). (Note: Use the general exponential function.)

OpenStudy (anonymous):

Start with y=Ca^x. Plug in the coordinate (-1,4), to get 4=Ca^-1. Solve for "a" to get, a=C/4. Substitute the known value of "a" back into the equation to get y=C(C/4)^x. Now plug in the coordinate (1,1) and solve for "C" to get C=2. Plug 2 into a=C/4 for "C" and solve to get a=1/2. Therefore y=2(1/2)^x will contain both of the points you provided